Question1524: Forty liters of a 60% salt solution are reduced to a 45% solution. How much solution must be drained off and replaced with distilled water so that the resulting solution contains only 45% solution.
Answer by richwmiller(17219)

I changed to viewpoint of water
0.4*40-0.4x+1x=0.55(40)
16+0.6x=22
0.6x=22-16
0.6x=6
x=10
check
0.4*30+1*10=0.55(40)
12+10=22
22=22
ok